# Affective Computing in the Era of Large Language Models: A Survey from the NLP Perspective

### v1

@@ 6.2 Affective Generation
This section introduces the benchmarks in the field of AG covering Emotional Support Conversation and Empathetic Response Generation.
@@ 6.2.1 Emotional support conversation
Emotional support conversation (ESC) aims to alleviate individuals’ emotional intensity and provide guidance for navigating personal challenges through engaging dialogue. The performance of LLMs on ESConv is initially analyzed, called LLMs2ESC [174]. LLMs2ESC finds that wide range of LLMs exhibits different preferences for support strategies and proposes a new suite of metrics that focus on strategies: proficiency, preference, and preference bias.
Cue-CoT [61] is a benchmark consisting of 6 in-depth dialogue datasets in both Chinese and English, considering three major aspects of user statuses: personality, emotions, and psychology, exhibited during the conversation, forming a comprehensive evaluation benchmark incorporating various user statuses in the context of dialogue response generation. It uses Automatic Evaluation and Human Evaluation.
Emotional Chain-of-Thought (ECoT) is a plug-and-play prompting method that enhances the performance of LLMs on various emotional generation tasks by aligning with human emotional intelligence guidelines [177]. Emotional Generation Score (EGS) is an automatic evaluation method to evaluate the quality of LLMs in emotional conditioned generation on a benchmark. This benchmark contains both textual and multimodal subsets. The textual benchmark contains five datasets: IEMOCAP [193], DailyDialog [65], EmpatheticDialogues [198], ESConv [12], and PENS [199]. The multimodal benchmark contains two datasets, SentiCap [200] and COCO [201].
FEEL [186] is a framework for evaluating Emotional Support Capability (ESC) using LLMs as evaluators. It constructs ESCEval, a high-quality dataset composed of human evaluations of emotional support capabilities, using instances from ESConv [12] and AUGESC [202], and trains the weights of LLMs on this dataset. The evaluation of ESC is categorized into two dimensions: Emotional Support Skills (Informativeness, Comprehensibility, Helpfulness) and Text Quality (Consistency, Coherence, Safety), using various metrics.
@@ 6.2.2 Empathetic Response Generation
MEDIC [187] is the first multimodal empathy dataset for psychotherapy scenarios considering visual, audio, and textual modalities. This paper proposes a new conceptual method for describing empathy, consisting of three mechanisms: Expression of Experience (EE), Emotional Reaction (ER), and Cognitive Reaction (CR).
To evaluate the empathy ability of LLMs, etc., how their feelings change when presented with specific situations. EmotionBench [188] collects 428 distinct situations, categorized into 36 factors, and conducts a comprehensive survey in the emotion appraisal theory. EmotionBench sets up a direction to align the emotional responses of LLMs with humans, focusing on eight negative emotions.
A comprehensive empirical investigation is conducted on the performance of LLMs, represented by gpt-3.5-turbo, in empathetic dialogue. For convenience, we refer to it as ChatGPT2EGR [189]. This study proposes three improvement methods: semantically similar in-context learning, two-stage interactive generation, and combination with a knowledge base. It adopts Distinct-n (Dist-1/2), BERTscore (PBERT, RBERT, FBERT), and BLEU-n (B2/4) as the main automatic metrics for evaluating response generation performance and samples 100 dialogues from the testing dataset for human evaluation.
STICKERCONV [55] is a multimodal empathetic dialogue dataset, comprising users with 2K generated personalities, 12.9K dialogue sessions, and 5.8K unique stickers. Unlike other benchmarks, STICKERCONV generates textual empathetic responses and outputs stickers as a visual modality to enhance empathy. It uses Automatic Evaluation, Human Evaluation, and LLM-based Evaluation.
@@ 6.2.3 Emotional Intelligence
In the context of large language models (LLMs), emotional intelligence (EI) has emerged as a critical area of focus. EI encompasses the capacity to identify, comprehend, regulate, and apply emotions across diverse situations [203]. To this end, a standardized EI test, Situational Evaluation of Complex Emotional Understanding (SECEU), suitable for both humans and LLMs is developed [190].
A novel benchmark, EQ-Bench [113], is designed to evaluate aspects of EI in Large Language Models (LLMs). EQ-Bench test dataset is constructed by GPT-4 to generate the dialogues to serve as the context for the test questions.
EmoBench [191] is a theory-based comprehensive EI benchmark for LLMs evaluation, consisting of a set of 400 hand-crafted questions, available in English and Chinese. The framework presents an extensive definition for machine’s EI, covering its essential capabilities: emotional understanding and emotional application.
EIBENCH [145] is a large curated collection of EI-related tasks converted into a text-to-text format, covering 15 tasks with 88 datasets. It primarily focuses on three aspects of the EI domain: Emotion Perception, Emotion Cognition, and Emotion Expression. To evaluate a model’s General Intelligence, EIBENCH conducts evaluations across four crucial dimensions: World Knowledge, General Reasoning, Commonsense Reasoning, and Reading Comprehension.
@@ 6.3 Others
This section introduces other benchmarks that do not easily fit into AU or AG tasks. For instance, one benchmark evaluates AU and AG, while another assesses multiple multimodal reasoning tasks with multimodal contexts, including multimodal sentiment analysis, sarcasm detection, and other tasks.
Executing a wide evaluation of the performance of InstructGPT and GPT-4, on a wide range of affective computing tasks, including aspect extraction, aspect polarity classification, opinion extraction, sentiment analysis, sentiment intensity detection, suicide tendency detection, toxicity detection, well-being assessment, engagement measurement, personality assessment and sarcasm detection, etc.. For convenience, we call this evaluation as ChatGPT2AC [99]. The automatic evaluation metrics are Unweighted Average Recall (UAR) and Accuracy.
MM-INSTRUCTEVAL [192] conducts comprehensive evaluations and establishs benchmarks for 31 models, including 23 MLLMs, using ten instructions across 16 diverse multimodal reasoning datasets with vision-text contexts. It uses automatic evaluation metrics such as Best Performance, Mean Relative Gain, Stability and Adaptability.
@@ 7 Discussion and Future Works
    Effective Evaluation
    Although LLMs have achieved excellent performance in various tasks, it remains essential to explore how to train an LLM that can effectively perceive sentiments and evaluate performance. When training LLMs dedicated to AC tasks, researchers face the challenge of not only retaining the original abilitie of modles, such as instruction understanding and reasoning, but also strengthening new abilities, specifically sentiment perception and cognition. Additionally, it is crucial to ensure the integrity of values, facts, and logic within the LLMs. In terms of evaluation, previous studies conducted before the emergence of LLMs proposed various metrics to measure model performance, such as Accuracy, BLEU, and ROUGE. However, these metrics are insufficient to fully assess the ability of LLMs, e.g., generalization, generation, perception, and cognitive. Establishing a comprehensive and effective evaluation standard remains an urgent problem to be solved.
@@
    Multimodal Affective Computing
    Large Language Models (LLMs) that focus solely on the text modality have been expanded to Multimodal Large Language Models (MLLMs) that can process both image and text modalities. In the field of AC, data from different modalities such as text, visual, audio, electroencephalogram (EEG), and electrooculogram (EOG) can express individual sentiments. Effectively collecting and utilizing this multimodal data to train unified MLLMs for affective computing tasks is crucial. Addressing challenges such as micro-expression recognition, audio/video sentiment capture, and advancing brain-computer interface and emotional intelligence technologies represents significant research directions.
@@
    Real-time Affective Computing
    The first step in sentiment analysis is building a dataset and performing sentiment analysis. Real-time sentiment analysis in a dynamic environment, by integrating data in real time and identifying and analyzing sentiment, is particularly critical for various practical applications such as online customer service, real-time monitoring, and interactive entertainment. Real-time sentiment analysis requires collecting and processing data in real-time, which presents several challenges—ensuring the real-time performance of data processing and sentiment analysis, especially maintaining time synchronization and consistency across different modalities. Real-time systems often encounter various environmental noises. Another major challenge is ensuring that the system can adapt to environmental changes and noise interference while maintaining the accuracy and stability of sentiment recognition.
@@
    Vertical Domain of Affective Computing LLMs
    LLMs perform well on most tasks. However, there is room for improvement in more challenging tasks, such as Joint Aspect-Based Sentiment Analysis and Emotion-Cause Pair Extraction. Models specifically tailored for affective computing within particular industry sectors or areas of expertise, such as depression detection, psychological counseling, and sentiment companionship, remain rare. Developing effective Vertical Domain LLMs for Affective Computing, capable of simultaneously handling both AU and AG tasks, is a valuable area of study.
@@ 8 Conclusion
Affective Computing (AC) has emerged as a crucial research direction in artificial intelligence. With the rise of Large Language Models (LLMs), significant progress has been made in understanding and generating emotions. This paper provides a comprehensive review of LLM applications in AC, exploring the roles of techniques such as instruction tuning and prompt engineering in affective understanding and generation tasks. We further conclude benchmarks for multiple large language models in AC to offer in-depth assistance to researchers and practitioners in related fields. Research indicates that LLMs, leveraging their powerful contextual learning and sequence generation capabilities, have demonstrated excellent affective understanding and generation performance. However, LLMs still face numerous challenges in multilingual and multicultural sentiment analysis, real-time affective computing, and emotion-related tasks in specific vertical domains. We discuss challenges and future research directions related to AC, and aim to provide insights into this rapidly advancing field.
@@ Acknowledgments
This research is supported by the National Natural Science Foundation of China (No.62272092, No.62172086, and No.62106039).

### v2

@@ 7.2 Affective Generation
This section introduces the benchmarks in the field of Affective Generation (AG), covering Emotional Support Conversation and Empathetic Response Generation.
@@ 7.2.1 Emotional Support Conversation
Emotional support conversation (ESC) aims to alleviate individuals’ emotional intensity and provide guidance for navigating personal challenges through engaging dialogue. Cue-CoT Wang et al. (2023a) is a benchmark consisting of six in-depth dialogue datasets in both Chinese and English, considering three aspects of user statuses: personality, emotions, and psychology. It forms a comprehensive evaluation benchmark for dialogue response generation, employing both Automatic Evaluation and Human Evaluation.
ESC-Eval Zhao et al. (2024a) is a framework for evaluating Emotion Support Conversation (ESC) in LLMs. It uses a specialized role-playing agent (ESC-Role) and a benchmark of 2,801 role cards to simulate multi-turn dialogues. Evaluation is performed via manual annotation across seven dimensions, such as fluency and empathy, enabling a dynamic assessment that moves beyond static, ground-truth-reliant metrics.
@@ 7.2.2 Empathetic Response Generation
MEDIC Zhu et al. (2023b) is a multimodal empathy dataset for psychotherapy scenarios, considering visual, audio, and textual modalities. It proposes a conceptual framework for empathy with three mechanisms: Expression of Experience (EE), Emotional Reaction (ER), and Cognitive Reaction (CR).
EmotionBench tse Huang et al. (2023) collects 428 distinct situations categorized into 36 factors and is designed to align LLM emotional responses with human appraisal, focusing on eight negative emotions.
STICKERCONV Zhang et al. (2024f) is a multimodal empathetic dialogue dataset with 2K generated personalities, 12.9K dialogue sessions, and 5.8K unique stickers. It generates textual empathetic responses and outputs stickers as a visual modality. Evaluation is conducted using Automatic Evaluation, Human Evaluation, and LLM-based Evaluation.
@@ 7.2.3 Emotional Intelligence
In the context of large language models (LLMs), emotional intelligence (EI) has emerged as a critical area of focus. EI encompasses the capacity to identify, comprehend, regulate, and apply emotions across diverse situations Mayer et al. (2016).
SECEU (Situational Evaluation of Complex Emotional Understanding) Wang et al. (2023d) is a standardized EI test suitable for both humans and LLMs. It evaluates the ability to comprehend and apply emotions in diverse situations.
EQ-Bench Paech (2024) is a benchmark designed to evaluate aspects of EI in LLMs. The test dataset is constructed using GPT-4 to generate dialogues that serve as the context for the test questions.
EmoBench Sabour et al. (2024) is a theory-based comprehensive EI benchmark for LLM evaluation, consisting of 400 hand-crafted questions available in English and Chinese. The framework defines machine EI in terms of emotional understanding and emotional application.
EIBENCH Zhao et al. (2024b) is a large collection of EI-related tasks in a text-to-text format, spanning 15 tasks across 88 datasets. It evaluates three key EI aspects (Emotion Perception, Cognition, and Expression) and general intelligence.
EmoVIT Xie et al. (2024) is the first instruction-tuned benchmark for visual emotion understanding, connecting visual cues with emotional cognition. It features a large-scale dataset of image-instruction pairs for tasks including classification, interpretation, and emotion-grounded conversation.
EMER Lian et al. (2023a) is a benchmark for Explainable Multimodal Emotion Recognition. It requires models to provide explanations for predicted emotions, evaluating their ability to integrate information from audio, video, and text modalities.
MER-UniBench Lian et al. (2025) is the largest benchmark for emotion recognition, with 115K human-verified samples across over 2K fine-grained categories. It supports tasks like fine-grained/basic emotion recognition and sentiment analysis.
@@ 7.3 Others
This section introduces other benchmarks that do not easily fit into AU or AG tasks, including affective computing and multimodal reasoning benchmarks.
ChatGPT2AC Amin et al. (2024) evaluates InstructGPT and GPT-4 across a wide range of affective computing tasks, including sentiment analysis, toxicity detection, and personality assessment. The benchmark uses Unweighted Average Recall (UAR) and Accuracy as its automatic evaluation metrics.
MM-INSTRUCTEVAL Yang et al. (2024b) establishes a benchmark for 31 models, including 23 MLLMs, using ten instructions across 16 diverse multimodal reasoning datasets with vision-text contexts. Evaluation metrics include Best Performance, Mean Relative Gain, Stability, and Adaptability.
@@ 8 Discussion and Future Works
    Multimodal Affective Computing
    Large Language Models (LLMs), initially designed to work exclusively with text, have now evolved into Multimodal Large Language Models (MLLMs), capable of processing both text and image modalities. In the domain of AC, data from diverse modalities—such as text, visuals, audio, electroencephalogram (EEG), and electrooculogram (EOG)—can convey individual emotions and sentiments. However, most current general-purpose MLLMs face limitations in their ability to integrate and interpret information from multiple modalities in a seamless manner. This often results in misinterpretations, particularly when critical emotional cues are absent or overlooked. To advance AC, it is essential to effectively collect and utilize multimodal data to train unified MLLMs. Key challenges include recognizing micro-expressions, capturing sentiment from audio and video, and enhancing brain-computer interface technologies and emotional intelligence systems. Addressing these challenges represents crucial research directions for the future.
@@
    Real-time Affective Computing
    The first step in sentiment analysis is building a dataset and performing sentiment analysis. Real-time sentiment analysis in a dynamic environment, by integrating data in real-time and identifying and analyzing sentiment, is particularly critical for various practical applications such as online customer service, real-time monitoring, and interactive entertainment. Real-time sentiment analysis requires collecting and processing data in real-time, which presents several challenges—ensuring the real-time performance of data processing and sentiment analysis, especially maintaining time synchronization and consistency across different modalities. Real-time systems often encounter various environmental noises. Another major challenge is ensuring that the system can adapt to environmental changes and noise interference while maintaining the accuracy and stability of sentiment recognition.
@@
    Reasoning in Affective Computing
    Current affective computing tasks primarily focus on understanding emotions and generating emotion-related content. However, they often overlook the deeper, underlying causes that trigger emotional states—specifically, the reasons behind the emergence, persistence, or fluctuation of certain emotions in response to various stimuli or contexts. Gaining insight into these triggers is essential for advancing systems beyond surface-level emotion recognition toward a more nuanced and comprehensive understanding of human affect. A major challenge lies in the extraction and integration of meaningful emotional cues to enable accurate sentiment reasoning. Successfully addressing this challenge could significantly enhance the interpretability and reliability of affective computing systems. However, evaluating and validating the sentiment reasoning remains a significant obstacle, requiring innovative approaches to balance technical accuracy with user comprehension.
@@
    Effective Evaluation
    Although LLMs have achieved excellent performance in various tasks, it remains essential to explore how to train an LLM that can effectively perceive sentiments and evaluate performance. When training LLMs dedicated to AC tasks, researchers face the challenge of not only retaining the original abilities of models, such as instruction understanding and reasoning, but also strengthening new abilities, specifically sentiment perception and cognition. Additionally, it is crucial to ensure the integrity of values, facts, and logic within the LLMs. In terms of evaluation, previous studies conducted before the emergence of LLMs proposed various metrics to measure model performance, such as Accuracy, BLEU, and ROUGE. However, these metrics are insufficient to fully assess the ability of LLMs, e.g., generalization, generation, perception, and cognitive. Establishing a comprehensive and effective evaluation standard remains an urgent problem to be solved.
@@ 9 Conclusion
Affective Computing (AC) has emerged as a crucial research direction in artificial intelligence. With the rise of Large Language Models (LLMs), significant progress has been made in understanding and generating emotions. This paper provides a comprehensive review of LLM applications in AC, exploring the roles of techniques such as instruction tuning, prompt engineering, and reinforcement learning in affective understanding and generation tasks. We further conclude benchmarks for multiple large language models in AC to offer in-depth assistance to researchers and practitioners in related fields. Research indicates that LLMs, leveraging their powerful contextual learning and sequence generation capabilities, have demonstrated excellent affective understanding and generation performance. However, LLMs still face numerous challenges in multilingual and multicultural sentiment analysis, real-time affective computing, and emotion-related tasks in specific vertical domains. We discuss challenges and future research directions related to AC, and aim to provide insights into this rapidly advancing field.
@@ Acknowledgments
Thanks to all co-authors for their hard work. The work is supported by the National Natural Science Foundation of China (No.62272092, No.62172086, and No.62106039), and the Fundamental Research Funds for the Central Universities under Grants (N25XQD004, N25ZLL045).